OpenStudy (anonymous):

Write the product as a sum or difference of trigonometric functions. 10 sin 56° cos 103°

OpenStudy (anonymous):

\[\sin x \cos y = \frac{1}{2}(\sin (x+y) + \sin(x - y))\]